# Magnetic properties of van der Waals layered single crystals DyOBr and SmOCl

### v1

@@ IV Conclusions
In summary, We have successfully synthesized the single crystals of two van der Waals magnetic insulators DyOBr and SmOCl. DyOBr is an antiferromagnet with a Néel temperature of 9.5 K and a band gap of 5.41 eV. It has a strong magnetic anisotropy along the -axis and susceptibility anomaly at T*=30 K due to the crystal field effect. Besides, DyOBr exhibits fractional magnetization plateau near 5/12 M, suggesting the existence of some field-induced quantum magnetic state. SmOCl is an antiferromagnet with a Néel temperature of 7.1 K and a band gap of 4.74 eV. In contrast, SmOCl exhibits strong Ising-like magnetic anisotropy with the easy axis to be -axis, which could be well understood by our crystal field calculations. Both compounds can be exfoliated mechanically into nanoflakes, demonstrating their potential applications in design vdW heterostructures and new-generation spintronic devices.

### v2

@@ IV Conclusions
In summary, we have successfully synthesized the single crystals of two van der Waals magnetic insulators DyOBr and SmOCl. DyOBr is an antiferromagnet with a Néel temperature of 9.5 K and a band gap of 5.41 eV. It has a strong magnetic anisotropy along the -axis and susceptibility anomaly at T*=30 K possibly due to the crystal field effect. Besides, DyOBr exhibits a one-third magnetization plateau, suggesting the existence of some exotic field-induced magnetic state. SmOCl is an antiferromagnet with Néel temperature of 7.1 K and a band gap of 4.74 eV. In contrast, SmOCl exhibits strong Ising-like magnetic anisotropy with the easy axis to be -axis, which could be well understood by our crystal field calculations. Both compounds can be exfoliated mechanically into nanoflakes, demonstrating their potential applications in design vdW heterostructures and new-generation spintronic devices.